# TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision

### v2

@@ Abstract
Hierarchical text classification aims to categorize each document into a set of classes in a label taxonomy. Most earlier works focus on fully or semi-supervised methods that require a large amount of human annotated data which is costly and time-consuming to acquire. To alleviate human efforts, in this paper, we work on hierarchical text classification with the minimal amount of supervision: using the sole class name of each node as the only supervision. Recently, large language models (LLM) show competitive performance on various tasks through zero-shot prompting, but this method performs poorly in the hierarchical setting, because it is ineffective to include the large and structured label space in a prompt. On the other hand, previous weakly-supervised hierarchical text classification methods only utilize the raw taxonomy skeleton and ignore the rich information hidden in the text corpus that can serve as additional class-indicative features. To tackle the above challenges, we propose TELEClass, Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification, which (1) automatically enriches the label taxonomy with class-indicative terms to facilitate classifier training and (2) utilizes LLMs for both data annotation and creation tailored for the hierarchical label space. Experiments show that TELEClass can outperform previous weakly-supervised methods and LLM-based zero-shot prompting methods on two public datasets.
TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision
Yunyi Zhang1, Ruozhen Yangthanks: Equal contribution1, Xueqiang Xu11footnotemark: 11, Rui Li11footnotemark: 12, Jinfeng Xiao1, Jiaming Shen3, Jiawei Han1, 1University of Illinois Urbana-Champaign, 2University of Science and Technology of China, 3Google DeepMind, yzhan238@illinois.edu
1 Introduction
Hierarchical text classification, aiming to classify documents into one or multiple classes in a label taxonomy, is an essential and fundamental task in text mining and natural language processing. Compared with standard text classification where label space is flat and relatively small (e.g., less than 20 classes), hierarchical text classification is more challenging given the larger and more structured label space and the existence of fine-grained and long-tail classes. Most earlier works tackle this task in fully supervised You et al. (2019); Huang et al. (2019); Chen et al. (2021) or semi-supervised settings Berthelot et al. (2019); Gururangan et al. (2019), and different models are proposed to learn from a substantial amount of human-labeled data. However, acquiring human annotation is often costly, time-consuming, and not scalable.
@@ 1 Introduction
Figure 1: An example document tagged with 3 classes. We automatically enrich each node with class-indicative terms and utilize LLMs to facilitate classification.
@@
Recently, large language models (LLM) such as GPT-4 OpenAI (2023) and Claude 3 Anthropic (2024) have demonstrated strong performance in flat text classification through zero-shot or few-shot prompting Sun et al. (2023). However, applying LLMs in hierarchical settings, with large and structured label spaces, remains challenging. Directly including hundreds of classes in prompts is ineffective, leading to structural information loss, increased computational costs, and diminished clarity for LLMs at distinguishing class-specific information. Along another line of research, Meng et al. (2019) propose to train a moderate-size text classifier by utilizing a small set of keywords or labeled documents for each class. However, compiling keyword lists for hundreds of classes and obtaining representative documents for each specific and niche category still demand significant human efforts. Shen et al. (2021) study the hierarchical text classification with minimal supervision, which takes the class name as the only supervision signal. Specifically, they introduce TaxoClass which generates pseudo labels with a textual entailment model for classifier training. However, this method overlooks additional class-relevant features in the text and suffers from the unreliable pseudo label selection due to domain shift of the entailment model.
@@
In this study, we advance minimally supervised hierarchical text classification by enriching the weak supervision with additional textual clues and leveraging LLMs tailored for the hierarchical label structure. First, we enrich the existing label taxonomy with class-specific terms derived from two sources: LLM generation and automated extraction from the corpus. For example, the “conditioner” class in Figure 1 is enriched with key terms like “moisture” and “soft hair”, which distinguish it from other classes. These terms enhance the supervision signal by combining knowledge from LLMs and text corpus and improve the pseudo label quality for classifier training. Second, we further explore the application of LLMs in hierarchical text classification from two perspectives. Specifically, we enhance LLM annotation efficiency and effectiveness through a taxonomy-guided candidate search and optimize LLM document generation to create more precise pseudo data by using taxonomy paths.
@@
Leveraging the above ideas, we introduce TELEClass: Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification. TELEClass consists of four major steps: (1) LLM-Enhanced Core Class Annotation, where we identify document “core classes” (i.e., fine-grained classes that most accurately describe the documents) by first enriching the taxonomy with LLM-generated key terms and then finding candidate classes with a top-down tree search algorithm for LLM to select the most precise core classes. (2) Corpus-Based Taxonomy Enrichment, where we analyze the taxonomy structure to additionally identify class-indicative topical terms through semantic and statistical analysis on the corpus. (3) Core Class Refinement with Enriched Taxonomy, where we embed documents and classes based on the enriched label taxonomy and refine the initially selected core classes by identifying the most similar classes for each document. (4) Text Classifier Training with Path-Based Data Augmentation, where we sample label paths from the taxonomy and guide the LLM to generate pseudo documents most accurately describing these fine-grained classes. Finally, we train the text classifier on two types of pseudo labels, the core classes and the generated data, with a simple text matching network and multi-label training strategy.
The contributions of this paper are summarized as follows: (1) We propose TELEClass, a new method for minimally supervised hierarchical text classification, which requires only the class names of the label taxonomy as supervision to train a multi-label text classifier. (2) We propose to enrich the label taxonomy with class-indicative terms, based on which we utilize an embedding-based document-class matching method to improve the pseudo label quality. (3) We study two ways of adopting large language models to help hierarchical text classifier training, which can improve the pseudo label quality and solve the data scarcity issue for fine-grained classes. (4) Extensive experiments on two datasets show that TELEClass can outperform strong weakly-supervised hierarchical text classification methods and zero-shot LLM prompting methods.
2 Problem Definition
@@
The minimally-supervised hierarchical text classification task aims to train a text classifier that can categorize each document into multiple nodes on a label taxonomy by using the name of each node as the only supervision Shen et al. (2021). For example, in Figure 1, the input document is classified as “hair care”, “shampoo”, and “scalp treatment”.
@@
Figure 2: Overview of the TELEClass framework.
We assume the label taxonomy to be a DAG instead of a tree, because it aligns better with real applications as one node can have multiple parents with different meanings. It is also more challenging because the classifier needs to assign a document multiple labels in different levels and paths.
3 Methodology
@@
3.1 LLM-Enhanced Core Class Annotation
Inspired by previous studies, we first tag each document with its “core classes”, which are defined as a set of classes that can describe the document most accurately Shen et al. (2021). This process also mimics the process of human performing hierarchical text classification: first select a set of most essential classes for the document and then trace back to their relevant classes to complete the labeling. For example, in Figure 1, by first tagging the document with “shampoo” and “scalp treatment”, we can easily find its complete set of classes.
@@ 3.1 LLM-Enhanced Core Class Annotation
$sim(c,d)=\max_{t\in T^{\text{LLM}}_{c}}\cos(\vec{t},\vec{d}),\vspace{-0.5em}$ (1)
where $\vec{\circ}$ denotes a vector representation by a pre-trained semantic encoder (e.g., Sentence Transformer Reimers and Gurevych (2019)).
@@
3.2 Corpus-Based Taxonomy Enrichment
@@
We first collect a set of relevant documents $D^{0}_{c}\subset\mathcal{D}$ for each class $c$, which contains all the documents whose initial core classes contain $c$ or its descendants. Then, inspired by Tao et al. (2018); Zhang et al. (2023b), we consider the following three factors for class-indicative term selection: (1) Popularity: a class-indicative term $t$ of a class $c$ should be frequently mentioned by its relevant documents, which is quantified by the document frequency $\log(1+df(t,D^{0}_{c}))$. (2) Distinctiveness: a class-indicative term $t$ for a class $c$ should be infrequent in its siblings, which is quantified as the softmax of $BM25(t,D^{0}_{c})$ relevance scores over the set of siblings, $Sib(c,c_{p})$. (3) Semantic similarity: a class-indicative term should also be semantically similar to the class name, which is quantified as the cosine similarity between their embeddings derived from a pre-trained encoder (e.g., BERT Devlin et al. (2019)). Finally, we define the affinity score between a term $t$ and a class $c$ corresponding to parent $p$ to be the geometric mean of the above scores, denoted as $\text{aff}(t,c,c_{p})$.
@@
To enrich the taxonomy, we first apply a phrase mining tool, AutoPhrase Shang et al. (2018), to mine quality single-token and multi-token phrases from the corpus as candidate terms [^2] Our method is flexible with any kinds of phrase mining methods like Gu et al. (2021).. Then, for each class $c$ and each of its parents $c_{p}$, we select the top-$k$ terms with the highest affinity scores with $c$ corresponding to $p$, denoted as $T(c,c_{p})$. Then, we take the union of these corpus-based terms together with the LLM-generated terms in the previous step to get the final enriched class-indicative terms for class $c$, $T_{c}=\left(\bigcup_{c_{p}\in Par(c)}T(c,c_{p})\right)\bigcup T^{\text{LLM}}_{c}$.
3.3 Core Class Refinement with Enriched Taxonomy
@@
With the enriched class-indicative terms for each class, we propose to further utilize them to refine the initial core classes. In this paper, we adopt an embedding-based document-class matching method. Unlike previous methods in flat text classification Wang et al. (2021a) that use keyword-level embeddings to estimate document and class representations, here, we are able to define class representations directly based on document-level embeddings thanks to the rough class assignments we created in the core class annotation step (c.f. Sec. 3.1).
@@
To obtain document representations, we utilize a pre-trained Sentence Transformer model Reimers and Gurevych (2019) to encode the entire document, which we denote as $\vec{d}$. Then, for each class $c$, we identify a subset of its assigned documents that explicitly mention at least one of the class-indicative keywords and thus most confidently belong to this class, $D_{c}=\{d\in D^{0}_{c}|\exists w\in T_{c},w\in d\}$. Then, we use the average of their document embeddings as the class representation, $\vec{c}=\frac{1}{|D_{c}|}\sum_{d\in D_{c}}\vec{d}$. Finally, we compute the document-class matching score as the cosine similarity between their representations.
@@
$\begin{split}conf_{i}=\text{diff}^{\,i}(m_{i}),\quad\mathbb{C}_{i}&=\{c^{i}_{1},\dots,c^{i}_{m_{i}}\},\\ m_{i}=\operatorname*{arg\,max}_{j\in\{1,\dots,|\mathcal{C}|-1\}}&\text{diff}^{\,i}(j).\end{split}$ (2)
Finally, we select the top 75% of documents $d_{i}$ and their refined core classes with the highest confidence scores $conf_{i}$, denoted as $\mathcal{D}^{\text{core}}$.
3.4 Text Classifier Training with Path-Based Data Augmentation
@@
Therefore, to overcome this issue, we propose the idea of path-based document generation by LLMs to generate a small number of augmented documents (e.g., $q=5$) for each distinct path from a level-1 node to a leaf node in the taxonomy. By adding the generated documents to the pseudo-labeled data, we can ensure that each class of the taxonomy will be a positive class of at least $q$ documents. Moreover, we use a path instead of a single class to guide the LLM generation, because the meaning of lower-level classes is often conditioned on their parents. For example, in Figure 2, a path “hair care” $\rightarrow$ “shampoo” can guide the LLM to generate text about hair shampoo instead of pet shampoo or carpet shampoo that are in different paths. To promote data diversity, we make one LLM query for each path and ask it to generate $q$ diverse documents. Appx. C shows the prompts we used. We denote the generated documents as $\mathcal{D}^{\text{gen}}$.
@@
Classifier architecture. We use a simple text matching network Shen et al. (2021) as our model architecture, which includes a document encoder initialized with a pre-trained BERT-base model Devlin et al. (2019) and a log-bilinear matching network. Class representations are initialized by class name embeddings (c.f. Sec. 3.2).
Training process. For each document with refined core classes, we construct its positive classes as the union of its refined core classes and their ancestors in the label taxonomy, and its negative classes are the ones that are not positive classes or descendants of any core class. This is because the ancestors of confident core classes are also likely to be true labels, and the descendants may not all be negative given that the automatically generated core classes are not optimal. For the LLM-generated documents, we are more confident in their labels, so we simply treat all the classes in the corresponding label path as positive classes and all other classes as negative.
We train the text-matching model with the standard binary cross-entropy loss. The loss terms of two sets of data are weighted by their relative size, $\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}$. Notice that we do not continue training the classifier with self-training that is commonly used in previous studies Meng et al. (2019); Shen et al. (2021). Using self-training may further improve the model performance, which we leave for future exploration. Algorithm 1 in Appendix summarizes TELEClass.
Table 1: Experiment results on Amazon-531 and DBPedia-298 datasets, evaluated by Example-F1, P@k, and MRR. The best score among zero-shot and weakly-supervised methods is boldfaced. “” indicates the numbers for these baselines are directly from previous paper Shen et al. (2021). “—” means the method cannot generate a ranking of predictions and thus MRR cannot be calculated.

### v3

@@ Abstract.
Hierarchical text classification aims to categorize each document into a set of classes in a label taxonomy, which is a fundamental web text mining task with broad applications such as web content analysis and semantic indexing. Most earlier works focus on fully or semi-supervised methods that require a large amount of human annotated data which is costly and time-consuming to acquire. To alleviate human efforts, in this paper, we work on hierarchical text classification with a minimal amount of supervision: using the sole class name of each node as the only supervision. Recently, large language models (LLM) have shown competitive performance on various tasks through zero-shot prompting, but this method performs poorly in the hierarchical setting because it is ineffective to include the large and structured label space in a prompt. On the other hand, previous weakly-supervised hierarchical text classification methods only utilize the raw taxonomy skeleton and ignore the rich information hidden in the text corpus that can serve as additional class-indicative features. To tackle the above challenges, we propose TELEClass, Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification, which combines the general knowledge of LLMs and task-specific features mined from an unlabeled corpus. TELEClass automatically enriches the raw taxonomy with class-indicative features for better label space understanding and utilizes novel LLM-based data annotation and generation methods specifically tailored for the hierarchical setting. Experiments show that TELEClass can significantly outperform previous baselines while achieving comparable performance to zero-shot prompting of LLMs with drastically less inference cost.
Weakly-Supervised Text Classification, Hierarchical Text Classification, Taxonomy Enrichment, Large Language Model
Equal Contribution.
journalyear: 2025copyright: ccconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714940isbn: 979-8-4007-1274-6/25/04ccs: Information systems Data miningccs: Computing methodologies Natural language processing
1. Introduction
Hierarchical text classification, aiming to classify documents into one or multiple classes in a label taxonomy, is a fundamental task in web text mining and NLP. Compared with standard text classification where label space is flat and relatively small (e.g., less than 20 classes), hierarchical text classification is more challenging given the larger and more structured label space and the existence of fine-grained and long-tail classes. Hierarchical text classification has broad applications such as web content organization (Dumais and Chen, 2000), semantic indexing (Li et al., 2019; Kang et al., 2024), and query classification (Cao et al., 2009; He et al., 2023, 2024b). Recent studies also show that hierarchically structured text (Sarthi et al., 2024; Edge et al., 2024) and document-level tagging (Poliakov and Shvai, 2024) can improve retrieval-augmented generation for large language models.
@@ 1. Introduction
Figure 1. An example document tagged with 3 classes. We automatically enrich each node with class-indicative terms and utilize LLMs to facilitate classification.
The key challenge of hierarchical text classification is how to understand the large structured label space to distinguish the semantics of similar classes. Most earlier works tackle this task in fully supervised (You et al., 2019; Huang et al., 2019; Chen et al., 2021) or semi-supervised settings (Berthelot et al., 2019; Gururangan et al., 2019), and different models are proposed to learn from a substantial amount of human-labeled data. However, acquiring human annotation is often costly, time-consuming, and not scalable.
Recently, large language models (LLM) such as GPT-4 (OpenAI, 2023) and Claude 3 (Anthropic, 2024) have demonstrated strong performance in flat text classification (Sun et al., 2023). However, applying LLMs in hierarchical settings remains challenging (Yu et al., 2023a). Directly including hundreds of classes in prompts is ineffective and inefficient, leading to structural information loss, diminished clarity for LLMs at distinguishing class-specific information, and prohibitively expensive inference cost given the long prompt for each test document.
@@
Along another line of research, Meng et al. (2019) propose to train a moderate-size text classifier by utilizing a small set of keywords or labeled documents for each class and a large unlabeled corpus. However, compiling keyword lists for hundreds of classes and obtaining representative documents for each specific and niche category still demand significant human efforts. Shen et al. (2021) study the hierarchical text classification with minimal supervision, which takes the class name as the only supervision signal. Specifically, they introduce TaxoClass which generates pseudo labels with a textual entailment model for classifier training. However, this method overlooks additional class-relevant features in the corpus that could be helpful for label space understanding. It also suffers from the unreliable pseudo label selection because the entailment model is not trained to compare which class is more relevant to the document.
@@
In this study, we advance minimally supervised hierarchical text classification by taking the advantage of both LLMs’ text understanding ability and task-specific knowledge of the unlabeled text corpus. First, we tackle the challenge of label space understanding by enriching the label taxonomy with class-specific terms derived from two sources: LLM generation and automated extraction from the corpus. For example, the “conditioner” class in Figure 1 is enriched with key terms like “moisture” and “soft hair”, which distinguish it from other classes. These terms enhance the supervision signal by combining knowledge from LLMs and text corpus and improve the pseudo label quality for classifier training. Second, we improve LLMs’ ability in hierarchical text classification from two perspectives: we enhance LLM annotation efficiency and effectiveness through a taxonomy-guided candidate search and also optimize LLM-based document generation to create more precise pseudo data by using taxonomy paths.
@@
Leveraging the above ideas, we introduce TELEClass: Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification. TELEClass consists of four major steps: (1) LLM-Enhanced Core Class Annotation, where we identify document “core classes” (i.e., fine-grained classes that most accurately describe the documents) by first enriching the taxonomy with LLM-generated key terms and then finding candidate classes with a top-down tree search algorithm for LLM to select the most precise core classes. (2) Corpus-Based Taxonomy Enrichment, where we analyze the taxonomy structure to additionally identify class-indicative topical terms through semantic and statistical analysis on the corpus. (3) Core Class Refinement with Enriched Taxonomy, where we embed documents and classes based on the enriched label taxonomy and refined the initially selected core classes by identifying the most similar classes for each document. (4) Text Classifier Training with Path-Based Data Augmentation, where we sample label paths from the taxonomy and guide the LLM to generate pseudo documents most accurately describing these fine-grained classes. Finally, we train the text classifier on two types of pseudo labels, the core classes and the generated data, with a simple text matching network and multi-label training strategy.
The contributions of this paper are summarized as follows:
  • We propose TELEClass, a new method for minimally supervised hierarchical text classification, which requires only the class names of the label taxonomy as supervision to train a multi-label text classifier.
  • We propose to enrich the label taxonomy with class-indicative terms, based on which we utilize an embedding-based document-class matching method to improve the pseudo label quality.
  • We study two ways of adopting large language models to hierarchical text classification, which can improve the pseudo label quality and solve the data scarcity issue for fine-grained classes.
  • Experiments on two datasets show that TELEClass can significantly outperform zero-shot and weakly-supervised hierarchical text classification baselines, while also achieving comparable performance to GPT-4 with drastically less inference cost.  [^1] Our code and datasets can be found at: https://github.com/yzhan238/TELEClass
2. Problem Definition
@@
The minimally-supervised hierarchical text classification task aims to train a text classifier that can categorize each document into multiple nodes on a label taxonomy by using the name of each node as the only supervision (Shen et al., 2021). For example, in Figure 1, the input document is classified as “hair care”, “shampoo”, and “scalp treatment”.
@@
Figure 2. Overview of the TELEClass framework.
We assume the label taxonomy to be a DAG instead of a tree, because it aligns better with real applications as one node can have multiple parents with different meanings. Therefore, the classifier needs to assign a document multiple labels in different levels and paths. Also, this setting does not restrict the prediction to be one label per level. In this way, the output can consist of labels from multiple paths or not necessarily contain a leaf node if the document is not specifically discussing a fine-grained topic. Such a setting is more flexible to fit various applications and also a more challenging research task.
3. Methodology
@@
3.1. LLM-Enhanced Core Class Annotation
Inspired by previous studies, we first tag each document with its “core classes”, which are defined as a set of classes that can describe the document most accurately (Shen et al., 2021). This process also mimics the process of human performing hierarchical text classification: first select a set of most essential classes for the document and then trace back to their relevant classes to complete the labeling. For example, in Figure 1, by first tagging the document with “shampoo” and “scalp treatment”, we can easily find its complete set of classes.
@@ 3.1. LLM-Enhanced Core Class Annotation
(1) $\small sim(c,d)=\max_{t\in T^{\text{LLM}}_{c}}\cos(\vec{t},\vec{d}),$
where $\vec{\circ}$ denotes a vector representation by a pre-trained semantic encoder (e.g., Sentence Transformer (Reimers and Gurevych, 2019)).
@@
3.2. Corpus-Based Taxonomy Enrichment
@@
We first collect a set of relevant documents $D^{0}_{c}\subset\mathcal{D}$ for each class $c$, which contains all the documents whose initial core classes contain $c$ or its descendants. Then, inspired by (Tao et al., 2018; Zhang et al., 2023c), we consider the following three factors for class-indicative term selection and adapt them to the hierarchical setting.
  • Popularity: a class-indicative term $t$ of a class $c$ should be frequently mentioned by its relevant documents, which is quantified by the log normalization of its document frequency,
    (2) $pop(t,c)=\log(1+df(t,D^{0}_{c})),$
    where $df(t,D)$ stands for the number of documents in $D$ that mention $t$.
  • Distinctiveness: a class-indicative term $t$ for a class $c$ should be infrequent in its siblings, which is quantified as the softmax of BM25 relevance function (Robertson and Zaragoza, 2009) over the set of siblings,
    (3) $dist(t,c,c_{p})=\frac{\exp(BM25(t,D^{0}_{c}))}{1+\sum_{c^{\prime}\in Sib(c,c_{p})}\exp(BM25(t,D^{0}_{c^{\prime}}))}.$
  • Semantic similarity: a class-indicative term $t$ should also be semantically similar to the class name of $c$, which is quantified as the cosine similarity between their embeddings derived from a pre-trained encoder (e.g., BERT (Devlin et al., 2019)), denoted as $sem(c,t)$.
Finally, we define the affinity score between a term $t$ and a class $c$ corresponding to parent $p$ to be the geometric mean of the above scores, denoted as $\text{aff}(t,c,c_{p})$.
@@
To enrich the taxonomy, we first apply a phrase mining tool, AutoPhrase (Shang et al., 2018), to mine quality single-token and multi-token phrases from the corpus as candidate terms [^3] Our method is flexible with any kinds of phrase mining methods like Gu et al. (2021).. Then, for each class $c$ and each of its parents $c_{p}$, we select the top-$k$ terms with the highest affinity scores with $c$ corresponding to $p$, denoted as $T(c,c_{p})$. Then, we take the union of these corpus-based terms together with the LLM-generated terms in the previous step to get the final enriched class-indicative terms for class $c$,
(4) $T_{c}=\left(\bigcup_{c_{p}\in Par(c)}T(c,c_{p})\right)\bigcup T^{\text{LLM}}_{c}.$
3.3. Core Class Refinement with Enriched Taxonomy
@@
With the enriched class-indicative terms for each class, we propose to further utilize them to refine the initial core classes. In this paper, we adopt an embedding-based document-class matching method. Unlike previous methods in flat text classification (Wang et al., 2021a) that use keyword-level embeddings to estimate document and class representations, here, we are able to define class representations directly based on document-level embeddings thanks to the rough class assignments we created in the core class annotation step (c.f. Sec. 3.1).
@@
To obtain document representations, we utilize a pre-trained Sentence Transformer model (Reimers and Gurevych, 2019) to encode the entire document, which we denote as $\vec{d}$. Then, for each class $c$, we identify a subset of its assigned documents that explicitly mention at least one of the class-indicative keywords and thus most confidently belong to this class, $D_{c}=\{d\in D^{0}_{c}|\exists w\in T_{c},w\in d\}$. Then, we use the average of their document embeddings as the class representation, $\vec{c}=\frac{1}{|D_{c}|}\sum_{d\in D_{c}}\vec{d}$. Finally, we compute the document-class matching score as the cosine similarity between their representations.
@@
(5) $\begin{split}conf_{i}=\text{diff}^{\,i}(m_{i}),\quad\mathbb{C}_{i}&=\{c^{i}_{1},\dots,c^{i}_{m_{i}}\},\\ m_{i}=\operatorname*{arg\,max}_{j\in\{1,\dots,|\mathcal{C}|-1\}}&\text{diff}^{\,i}(j).\end{split}$
Finally, we select top 75% of documents $d_{i}$ and their refined core classes with the highest confidence scores $conf_{i}$, denoted as $\mathcal{D}^{\text{core}}$.
3.4. Text Classifier Training with Path-Based Data Augmentation
@@
Therefore, to overcome this issue, we propose the idea of path-based document generation by LLMs to generate a small number of augmented documents (e.g., $q=5$) for each distinct path from a level-1 node to a leaf node in the taxonomy. By adding the generated documents to the pseudo-labeled data, we can ensure that each class of the taxonomy will be a positive class of at least $q$ documents. Because we generate a small constant number of documents for each label path, it also does not affect the distribution of the frequent classes. Moreover, we use a path instead of a single class to guide the LLM generation, because the meaning of lower-level classes is often conditioned on their parents. For example, in Figure 2, a path “hair care” $\rightarrow$ “shampoo” can guide the LLM to generate text about hair shampoo instead of pet shampoo or carpet shampoo that are in different paths. To promote data diversity, we make one LLM query for each path and ask it to generate $q$ diverse documents. We denote the generated documents as $\mathcal{D}^{\text{gen}}$. Appx. B shows the prompts we used.
@@
Classifier architecture. We use a simple text matching network similar to (Shen et al., 2021) as our model architecture, which includes a document encoder initialized with a pre-trained BERT-base model (Devlin et al., 2019) and a log-bilinear matching network. Class representations are initialized by class name embeddings (c.f. Sec. 3.2) and are detached from the encoder model, so only the embeddings will be updated without back-propagation to the backbone model. Formally, the classifier predicts the probability of document $d_{i}$ belonging to class $c_{j}$:
$\small p(c_{j}|d_{i})=\mathcal{P}(y_{j}=1|d_{i})=\sigma(\exp{(\mathbf{c}_{j}^{T}\mathbf{W}\mathbf{d}_{i})}),$
where $\sigma$ is the sigmoid function, $\mathbf{W}$ is a learnable interaction matrix, and $\mathbf{c}_{j}$ and $\mathbf{d}_{i}$ are the encoded class and document.
Training process. For each document $d_{i}$ tagged with core classes $\mathbb{C}_{i}$, we construct its positive classes $\mathbb{C}^{\text{core}}_{i,+}$ as the union of its core classes and their ancestors in the label taxonomy, and its negative classes $\mathbb{C}^{\text{core}}_{i,-}$ are the ones that are not positive classes or descendants of any core class. This is because the ancestors of confident core classes are also likely to be true labels, and the descendants may not all be negative given that the automatically generated core classes are not optimal. Formally,
$\small\mathbb{C}^{\text{core}}_{i,+}$ $=\mathbb{C}_{i}\cup\left(\cup_{c\in\mathbb{C}_{i}}Anc(c)\right),$
$\mathbb{C}^{\text{core}}_{i,-}$ $=\mathcal{C}-\mathbb{C}^{\text{core}}_{i,+}-\cup_{c\in\mathbb{C}_{i}}Des(c),$
where $Anc(c)$ and $Des(c)$ denote the set of ancestors and descendants and class $c$, respectively. For the LLM-generated documents, we are more confident in its pseudo labels, so we simply treat all the classes in the corresponding path as positive classes and all other classes as negative,
$\small\mathbb{C}^{\text{gen}}_{p,+}=\mathbb{C}_{p},\quad\mathbb{C}^{\text{gen}}_{p,-}=\mathcal{C}-\mathbb{C}_{p}.$
Then, we train a multi-label classifier with the binary cross entropy loss (BCE):
$\mathcal{L}^{\text{core}}=-\sum_{d_{i}\in\mathcal{D}^{\text{core}}}\left(\sum_{c_{j}\in\mathbb{C}^{\text{core}}_{i,+}}\log{p(c_{j}|d_{i})}+\sum_{c_{j}\in\mathbb{C}^{\text{core}}_{i,-}}\log{(1-p(c_{j}|d_{i}))}\right)$
$\mathcal{L}^{\text{gen}}=-\sum_{d^{p}_{i}\in\mathcal{D}^{\text{gen}}}\left(\sum_{c_{j}\in\mathbb{C}^{\text{gen}}_{p,+}}\log{p(c_{j}|d^{p}_{i})}+\sum_{c_{j}\in\mathbb{C}^{\text{gen}}_{p,-}}\log{(1-p(c_{j}|d^{p}_{i}))}\right)$
(6) $\small\mathcal{L}=\mathcal{L}^{\text{core}}+\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}\cdot\mathcal{L}^{\text{gen}}$
The loss terms of two sets of data are weighted by their relative size, $\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}$. Notice that we do not continue training the classifier with self-training that is commonly used in previous studies (Meng et al., 2019; Shen et al., 2021). Using self-training may further improve the model performance, which we leave for future exploration. Algorithm 1 in Appendix summarizes TELEClass.
Table 1. Datasets overview.
Dataset # unlabeled train # test # labels
Amazon-531 29,487 19,685 531
DBPedia-298 196,665 49,167 298
Table 2. Experiment results on Amazon-531 and DBPedia-298 datasets, evaluated by Example-F1, P@k, and MRR. The best score among zero-shot and weakly-supervised methods is boldfaced. “” indicates the numbers for these baselines are directly from previous paper (Shen et al., 2021). “—” means the method cannot generate a ranking of predictions and thus MRR cannot be calculated.